OpenStudy (ledah):

What will the exponent be when you simplify the following problem? (y^4)^2=?

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large ({a^x})^y = {a^{xy}} }\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow {(y^4)^2} = y^{4 \times 2} }\)

OpenStudy (anonymous):

\[(y^4)^2=y^(4*2)=y^8\] Also what Parth said